OpenStudy (cookiemonster2412):

Find the geometric mean of 5 and 8. If necessary, give the answer in simplest radical form.

OpenStudy (acxbox22):

multiply the 2 numbers then find the square root of the result

OpenStudy (acxbox22):

5 x 8=40 \[\sqrt{40}=2\sqrt{10}\]

OpenStudy (cookiemonster2412):



OpenStudy (acxbox22):

do you know how to simplify square roots?

OpenStudy (cookiemonster2412):

no

OpenStudy (acxbox22):

okay then, to simply a square root you have to look for any square factors like 4,9,16,25 etc for example the \[\sqrt{40}\] it has one square factor, 4 4 x 10 =40 and the square root of 4 is 2 so it is written as \[2\sqrt{10}\] the square root of the 4 goes outside and what remains stays under the radical

OpenStudy (cookiemonster2412):

oh ok thank you again

OpenStudy (acxbox22):

np